1. ABC Corp. has an ROE (return on reinvested earnings) of 20% and a dividend payout ratio of 40%. The next annual earnings are expected to be $3 per share (that is, EPS in year 1 is $3.00). The firm's required return on the stock is 17%. The value of the stock today is $____________.

2. Company A  just paid a $1.00 dividend per share and its future dividends are expected to grow at an annual rate of 6% for the foreseeable future. The beta of company A's stock is 1.25, the risk-free rate of return is 4% and the expected return on the market portfolio is 10.4%. The value of the stock today is $___________.
Answer #1

1. The value of the stock would be the present value of future dividend payments. The next earnings is expected at USD 3 per share. The dividend payout would be USD 3*0.4 = USD 1.2as dividend payout ratio is 40%

Therefore the present value of the stock would be a perpetuity whose value would be given by D1/r = 1.2/.017 = USD 7.05 per share

2. We first need to calculate the required rate of return on the stock using the CAPM (capital asset pricing model)

r = 4 + 1.25*(10.4-4) = 12% (as r = RFR + Beta*(Expected return on Market - RFR))

The present value of the stock as per dividend discount model would now be D1/r-g = 1*1.06/(.12-.06) = USD 17.67